1. The Excess Capacity theorem asserts that because a monopolistic competitor produces an output that is smaller than what it would to minimize costs of production, and therefore has

A. no right to export its good.

B excess capacity.

C a tendency to charge a price that is less than average total cost.

D an incentive to lobby for lower import tariffs.

2. Consider the production possibility curves for England and Portugal. Each country can produce cloth or wine. England can produce at most 10 units of cloth or at most 10 units of wine. Portugal can produce at most 4 units of cloth or 8 units of wine. Which country has a comparative advantage in the production of wine?

A. England

B Portugal

3. An autarkic country has a production possibility curve with constant opportunity costs such that it can produce at most 210 units of silk, or at most 70 units of kerosene. Suppose this country is currently producing 56 units of kerosene and 42 units of silk, and decides instead to produce 69 units of silk; how much kerosene will it produce?
Answer #1
  1. The correct option is (B). When the firm produces at a level less than the minimum point this means that the firm has excess capacity.
  2. Opportunity cost of cloth in England = 10/10 = 1

Opportunity cost of wine in England = 10/10 = 1

Opportunity cost of cloth in Portugal = 8/4 = 2

Opportunity cost of wine in Portugal = 4/8 = 0.5

Thus from comparing opportunity cost. England has high opportunity cost of Wine as compared to Portugal thus England has comparative advantage in production of wine.

Opportunity cost of cloth I greater in Portugal than England. Thus Portugal has comparative advantage in production of cloth.

The correct option is A. England

  1. Opportunity cost of silk in terms of kerosene = 70/210 = 1/3 unit of kerosene for one unit addition of silk

Change in silk production = 69 – 42 = 27 units

Thus change in kerosene required = -1/3*27 =-9 unit of kerosene.

New kerosene = 56 – 9 = 47 units of kerosene to be produced.